OpenStudy (anonymous):

Tom recorded the outdoor temperature for 8 consececutive days.Seven of those temperatures were 71,80,69,80,73,77,and 70.The average for all 8 days was 75.What was the temperature on the eighth day?

OpenStudy (anonymous):

\[\frac{ 71 + 80 + 69 + 80 + 73 + 77 + 70 + x }{ 8 } = 75\]This is taking the sum of all 8 days (x is the eighth day) and dividing that sum by 8 to get 75. So, you multiply both sides by 8 to start with.

OpenStudy (anonymous):

That will give you: 71 + 80 + 69 + 80 + 73 + 77 + 70 + x = (75)(8) So, just subtract the 7 recorded temperatures from each side to isolate "x".

OpenStudy (anonymous):

So, to sum up, you take the average of all 8 days by adding up all the values (where "x" is one of the values) and set that equal to the given average.

OpenStudy (anonymous):

Did you take 75x8

OpenStudy (anonymous):

That's the right-hand side

OpenStudy (anonymous):

Choose from one of these:72,75,78,80

OpenStudy (anonymous):

This one is hard

OpenStudy (anonymous):

Yes, your answer is among those 4. 71 + 80 + 69 + 80 + 73 + 77 + 70 + x = (75)(8) (71 + 80 + 69 + 80 + 73 + 77 + 70) + x = 600 x = 600 - (71 + 80 + 69 + 80 + 73 + 77 + 70) x = 600 - 520 x = 80

OpenStudy (anonymous):

Making sense yet?

OpenStudy (anonymous):

I guess if you say so & thank you

OpenStudy (anonymous):

lol You're welcome, but it's not too ok if it makes sense only to me! We can go over any part of it if you want.